# Large Language Models Are Human-Like Internally

### v1

@@ Abstract
Recent cognitive modeling studies have reported that larger language models (LMs) exhibit a poorer fit to human reading behavior Oh and Schuler (2023b); Shain et al. (2024); Kuribayashi et al. (2024), leading to claims of their cognitive implausibility. In this paper, we revisit this argument through the lens of mechanistic interpretability and argue that prior conclusions were skewed by an exclusive focus on the final layers of LMs. Our analysis reveals that next-word probabilities derived from internal layers of larger LMs align with human sentence processing data as well as, or better than, those from smaller LMs. This alignment holds consistently across behavioral (self-paced reading times, gaze durations, MAZE task processing times) and neurophysiological (N400 brain potentials) measures, challenging earlier mixed results and suggesting that the cognitive plausibility of larger LMs has been underestimated. Furthermore, we first identify an intriguing relationship between LM layers and human measures: earlier layers correspond more closely with fast gaze durations, while later layers better align with relatively slower signals such as N400 potentials and MAZE processing times. Our work opens new avenues for interdisciplinary research at the intersection of mechanistic interpretability and cognitive modeling.
Figure 1: Different measures of human sentence processing align with surprisal from different layers of language models (LMs), and the best layer is typically not the final one. Larger LMs can better simulate human reading data with their internal layer than smaller LMs.
@@ 1 Introduction
Previous studies have provided substantial evidence supporting expectation-based accounts of human sentence processing (Shain et al. 2024; inter alia). However, they reveal an intriguing trend: surprisal estimates from large language models (LLMs) often deviate from human reading behavior, and rather smaller models, such as GPT-2 small, offer better simulations of human behavior Shain et al. (2024); Oh and Schuler (2023b); Kuribayashi et al. (2022, 2024). This observation – larger LMs are less human-like — has sparked intriguing linguistic questions Wilcox et al. (2024) as well as a fair amount of confusion within the community. Why do smaller LMs appear more human-like, despite their generally poorer linguistic competence Waldis et al. (2024)?
In this work, we highlight the cognitively plausible aspects of LLMs, challenging existing conclusions. Specifically, we show that surprisal derived from the internal layers of larger LMs aligns with human sentence processing data as well as, or even better than, that from smaller LMs . Previous studies, focusing exclusively on final-layers’ surprisal, have overlooked this critical insight. These results could be drawn with techniques from mechanistic interpretability Dar et al. (2023); Belrose et al. (2023); Wendler et al. (2024), logit lens nostalgebraist (2020) or dubbed early exits Kaya et al. (2019); we compute next-word surprisals directly from internal layers of LMs by projecting intermediate representations into the output vocabulary space, bypassing subsequent layers. We additionally revealed that surprisal from earlier layers fits better with fast human responses (first-pass gaze durations and self-paced reading time), while surprisal from later layers aligns more closely with slower measures (N400 and MAZE task data) (Figure 1). This also resolves the previously suggested behavior–neurophysiology gap in LM-based cognitive modeling: smaller LMs predict reading behavior better Oh and Schuler (2023b), while larger LMs excel in modeling neurophysiological data Schrimpf et al. (2021); Michaelov et al. (2024a); Hosseini et al. (2024). We would suggest that this gap stemmed from the inconsistent treatment of internal layers (e.g., exclusive reliance on final layers, or inconsistent inclusion of intermediate layers). If all the internal layers were focused on, even larger LMs have human-like layers through the lens of both behavioral and neurophysiological data.
@@ 2.1 Cognitive modeling and NLP
A key objective in linguistics is to understand how humans process language Crocker (2007), a goal that remains pertinent even in the era of LLMs. According to perspectives outlined by Marr (1982), information processing can be examined at three levels: (i) the computational level—what is the goal of computation?, (ii) the algorithmic level—how does the model achieve the goal?, and (iii) the implementational level—how is it physically implemented?. Humans can be viewed as an information processing model, and surprisal theory Hale (2001); Levy (2008); Smith and Levy (2013) provides empirical evidence as a computational-level explanation of human sentence processing (Smith and Levy 2013; Frank et al. 2015; Shain et al. 2024; inter alia). [^1] Surprisal theory has also been critiqued van Schijndel and Linzen (2021); Huang et al. (2024), particularly for its failure to account for the cognitive load incurred in complex sentences. Our contribution will be orthogonal to such criticism (see Limitation section). According to surprisal theory, humans continuously predict upcoming information during reading, with cognitive load incurred by unpredictable information.
Once the computational-level theory is assumed, the next question can be more relevant to the algorithmic level: with what kind of algorithms and representations, humans predict upcoming information. The NLP community has developed various methods to compute next-word probabilities, ranging from N-gram models to LLMs. By exploring which NLP models best simulate human sentence processing, researchers can bridge the models demanded in psycholinguistics (language processing in humans) and those implemented in NLP (language processing in machines). This alignment is typically investigated by analyzing which LMs compute surprisal $-\log p(\mathrm{word}|\mathrm{context})$ that correlates with human measures (e.g., word-by-word gaze durations) based on the surprisal theory. Orthogonal to the exploration of various model architectures, this study investigates in which part/layer of a model human-like measures emerge.
@@ 2.2 Poorer fit of larger LMs’ surprisal to human reading behavior
Before the advent of LLMs, models such as incremental parsers Levy (2008), N-gram LMs Smith and Levy (2013), RNN-based LMs Frank and Bod (2011); Aurnhammer and Frank (2019), Transformer-based LMs Merkx and Frank (2021), and syntactic LMs Hale et al. (2018); Yoshida et al. (2021); Oh et al. (2021, 2022) were extensively analyzed in cognitive modeling. In the days of much smaller LMs than modern LLMs, LM-scaling generally improved their ability to simulate human sentence processing data Frank and Bod (2011); Goodkind and Bicknell (2018); Wilcox et al. (2020, 2023a). However, recent studies have questioned the generality of this scaling effect. The reversed trend was first found in typologically distant languages Kuribayashi et al. (2021), and even within English, further scaling up LMs have shown weaker alignment with human reading behavior Kuribayashi et al. (2022); Shain et al. (2024); Oh and Schuler (2023b). This bigger is not always better phenomenon has become a key focus of LM-based cognitive modeling Wilcox et al. (2024), with researchers investigating why LLMs appear cognitively implausible Kuribayashi et al. (2022); Oh and Schuler (2023a, b); Oh et al. (2024); Nair and Resnik (2023); Kuribayashi et al. (2024). This study contributes to this discussion from a different perspective: We show the cognitively plausible properties of LLMs rather than solely identifying their limitations.
In addition, recent studies have typically reported behavior–neurophysiology gap in LM-scaling effects Michaelov et al. (2024a); Aw et al. (2024); smaller LMs simulate reading behavior better Oh and Schuler (2023b), while larger LMs simulate neurophysiological data better Schrimpf et al. (2021); Michaelov et al. (2024a); Hosseini et al. (2024). Opposite effect of instruction-tuning was also observed between brain data and behavioral data Aw et al. (2024); Kuribayashi et al. (2024). We begin by offering a perspective to address this gap, showing that the internal layers of larger LMs are more effective at modeling both behavioral and neurophysiological data.
2.3 Mapping different human measures with different LM layers
@@
Furthermore, self-paced reading and eye-tracking measures often exhibit spillover effects, where the processing of one word influences subsequent words Rayner (1998). This suggests that the comprehension of a word extends beyond the immediate moment, and the associated reading times may only capture an early stage of processing. In contrast, the MAZE task Forster et al. (2009); Boyce and Levy (2023), which measures the time taken to select a plausible continuation from two candidates, mitigates spillover effects and is thus expected to reflect the full process of word processing. Our findings align with this perspective. Early LM layers are more closely aligned with gaze durations and self-paced reading times, while later layers show a stronger alignment with MAZE, as discussed in § 4.
Note that similar layer-wise LM-brain alignments have been attempted in fMRI modeling research Toneva and Wehbe (2019); Schrimpf et al. (2021); Caucheteux et al. (2023) and suggested that different brain areas better align with different LM layers. They typically trained linear regression models to predict brain activity directly using $d$-dimentional LM internal representations $\bm{h}\in\mathbb{R}^{d}$ as features, instead of using surprisal measures $-\log p(\mathrm{word}|\mathrm{context})\in\mathbb{R}_{\geq 0}$. Thus, their results are not comparable with exitsing surprisal-based studies and may rather suffer from a confounding factor of different $d$ for different LMs when precisely discussing LM-scaling effects.
2.4 Early exits of neural models’ prediction
In the engineering context, predictions from internal layers of deep neural networks (i.e., early exists of the results from internal layers) are used to improve inference efficiency by avoiding their overthiking Graves (2016); Banino et al. (2021), which is also called adaptive computation time Kaya et al. (2019); Zhou et al. (2020). Such technique has also recently been employed to enhance interpretability research to identify at which layer a particular prediction shapes nostalgebraist (2020); Dar et al. (2023); Belrose et al. (2023). We apply logit lens nostalgebraist (2020) and its variant of tuned lens Belrose et al. (2023), projecting intermediate representations to the output space, to extract next-word predictions from the internal layers of LLMs.
@@ 3.1 Human Data
Table 1 lists the 15 human reading datasets used in our study (we additionally use MECO in § 5.4), which include human measurements from various methods: self-paced reading time (SPR), first-pass gaze duration (FPGD), Maze task processing time (MAZE), and electroencephalography (EEG; specifically the N400 component). The datasets share a common format: each word $w_{t}$ is annotated with $\mathrm{Cost}(w_{t})\in\mathbb{R}$ representing the human cognitive load associated with it. Our corpus selection aligns with recent studies Kuribayashi et al. (2024); Michaelov et al. (2024a); de Varda et al. (2024); McCurdy and Hahn (2024). [^2] We applied the same preprocessing as Kuribayashi et al. (2024) (DC, NS), de Varda et al. (2024) (UCL), Hahn et al. (2022) (Fillers), and Michaelov et al. (2024a) (N400). For ZuCO, we only used the naturalistic reading part, and for its N400, we averaged the values at the central electrode between 300-500ms during the first pass over a word.

### v2

@@ Abstract
Recent cognitive modeling studies have reported that larger language models (LMs) exhibit a poorer fit to human reading behavior Oh and Schuler (2023b); Shain et al. (2024); Kuribayashi et al. (2024), leading to claims of their cognitive implausibility. In this paper, we revisit this argument through the lens of mechanistic interpretability and argue that prior conclusions were skewed by an exclusive focus on the final layers of LMs. Our analysis reveals that next-word probabilities derived from internal layers of larger LMs align with human sentence processing data as well as, or better than, those from smaller LMs. This alignment holds consistently across behavioral (self-paced reading times, gaze durations, MAZE task processing times) and neurophysiological (N400 brain potentials) measures, challenging earlier mixed results and suggesting that the cognitive plausibility of larger LMs has been underestimated. Furthermore, we first identify an intriguing relationship between LM layers and human measures: earlier layers correspond more closely with fast gaze durations, while later layers better align with relatively slower signals such as N400 potentials and MAZE processing times. Our work opens new avenues for interdisciplinary research at the intersection of mechanistic interpretability and cognitive modeling. [^1] Code is available at https://github.com/kuribayashi4/surprisal_internal_layers
@@ 1 Introduction
Previous studies have provided substantial evidence supporting expectation-based accounts of human sentence processing (Shain et al. 2024; inter alia). However, they reveal an intriguing trend: surprisal estimates from large language models (LLMs) often deviate from human reading behavior, and rather smaller models, such as GPT-2 small, offer better simulations of human behavior Shain et al. (2024); Oh and Schuler (2023b); Kuribayashi et al. (2022, 2024). This observation — larger LMs are less human-like — has sparked intriguing linguistic questions Wilcox et al. (2024) as well as a fair amount of confusion within the community. Why do smaller LMs appear more human-like, despite their generally poorer linguistic competence Waldis et al. (2024)?
In this work, we highlight the cognitively plausible aspects of LLMs, challenging existing conclusions. Specifically, we show that surprisal derived from the internal layers of larger LMs aligns with human sentence processing data as well as, or even better than, that from smaller LMs. Previous studies, focusing exclusively on final-layers’ surprisal, have overlooked this critical insight. These results could be drawn with techniques from mechanistic interpretability Dar et al. (2023); Belrose et al. (2023); Wendler et al. (2024), logit lens nostalgebraist (2020) or dubbed early exits Kaya et al. (2019); we compute next-word surprisals directly from internal layers of LMs by projecting intermediate representations into the output vocabulary space, bypassing subsequent layers. We additionally reveal that surprisal from earlier layers fits better with fast human responses (first-pass gaze durations and self-paced reading time), while surprisal from later layers aligns more closely with slower measures (N400 and MAZE task data) (Figure 1). This also resolves the previously suggested behavior–neurophysiology gap in LM-based cognitive modeling: smaller LMs predict reading behavior better Oh and Schuler (2023b), while larger LMs excel in modeling neurophysiological data Schrimpf et al. (2021); Michaelov et al. (2024a); Hosseini et al. (2024). We suggest that this gap stems from the inconsistent treatment of internal layers (e.g., exclusive reliance on final layers, or inconsistent inclusion of intermediate layers). If all the internal layers are focused on, even larger LMs have human-like responses through the lens of both behavioral and neurophysiological data.
Figure 1: Different measures of human sentence processing align with surprisal from different layers of language models (LMs), and the best layer is typically not the final layer. In the bottom plot, for example, gaze duration (blue dots) and EEG signal (red dots) correlate with earlier and later layers of LMs, respectively. Each dot corresponds to the fit of surprisal (y-axis) from a particular layer depth (x-axis) to human data (ZuCO corpus).
@@ 2.1 Cognitive modeling and NLP
A key objective in linguistics is to understand how humans process language Crocker (2007), a goal that remains pertinent even in the era of LLMs. According to perspectives outlined by Marr (1982), information processing can be examined at three levels: (i) the computational level: what is the goal of computation?; (ii) the algorithmic level: how does the model achieve the goal?; and (iii) the implementational level: how is it physically implemented?. Humans can be viewed as an information processing model, and surprisal theory — humans continuously predict upcoming information during reading, with cognitive load incurred by unpredictable information — Hale (2001); Levy (2008); Smith and Levy (2013) has accumulated its empirical evidence (Smith and Levy 2013; Frank et al. 2015; Shain et al. 2024; inter alia). [^2] Surprisal theory has also been critiqued van Schijndel and Linzen (2021); Huang et al. (2024), particularly for its failure to account for the cognitive load incurred in complex sentences. Our contribution is orthogonal to such criticism (see the Limitations section). An orthogonal, algorithmic-level question regarding the surprisal theory is with what kind of algorithms and representations, humans predict upcoming information. The NLP community has developed various methods to compute next-word probabilities, ranging from incremental parsers to LLMs, and researchers have tested their alignment with human reading data to provide insights into that question. This alignment is typically investigated by analyzing which LMs compute surprisal $-\log p(\mathrm{word}|\mathrm{context})$ that correlates with human measures (e.g., word-by-word gaze durations) based on the surprisal theory.
@@ 2.2 Poorer fit of larger LMs’ surprisal to human reading behavior
In the days of much smaller LMs/parsers than modern LLMs (Hale 2001; Levy 2008; Smith and Levy 2013; Frank and Bod 2011; Aurnhammer and Frank 2019; Merkx and Frank 2021; inter alia.), model-scaling generally improved their ability to simulate human sentence processing data Frank and Bod (2011); Goodkind and Bicknell (2018); Wilcox et al. (2020, 2023a). However, recent studies have questioned the generality of this scaling effect. The reversed trend was first found in typologically distant languages Kuribayashi et al. (2021), and even within English, further scaling up LMs have shown weaker alignment with human reading behavior Kuribayashi et al. (2022); Shain et al. (2024); Oh and Schuler (2023b). This bigger is not always better phenomenon has become a key focus of LM-based cognitive modeling Wilcox et al. (2024), with researchers investigating why LLMs appear cognitively implausible Kuribayashi et al. (2022); Oh and Schuler (2023a, b); Oh et al. (2024); Nair and Resnik (2023); Kuribayashi et al. (2024). In addition, from a more interdisciplinary view, mixed results are reported regarding such scaling effects. For example, smaller LMs simulate reading behavior better Oh and Schuler (2023b), while larger LMs simulate neurophysiological data better Schrimpf et al. (2021); Michaelov et al. (2024a); Hosseini et al. (2024). Opposite effect of instruction-tuning was also observed between brain data and behavioral data Aw et al. (2024); Kuribayashi et al. (2024). We offer a perspective to address these negative scaling effects and behavior–neurophysiology gap, showing that the internal layers of larger LMs are more effective at modeling both behavioral and neurophysiological data.
2.3 Human measure differences
@@
3 Methods
3.1 Probabilities from internal layers
Our main proposal is to use the next-word probability $p(w_{t}|\bm{w}_{<t})$ of word $w_{t}$ in its context $\bm{w}_{<t}=[w_{1},\cdots,w_{t-1}]^{\top}$ from internal layers in cognitive modeling, so we begin with how to extract internal surprisal. We use two methods of logit-lens nostalgebraist (2020) and its sophisticated version of tuned-lens Belrose et al. (2023). The first method, logit-lens, extracts the probability of a word $w_{t}$ from a $d$-dimentaional internal representation $\bm{h}_{l,t}\in\mathbb{R}^{d}$ at the $l$-th layer and time step $t$, as follows:
$=\mathrm{softmax}(\bm{W}_{U}\mathrm{LayerNorm}(\bm{h}_{l,t}))[\mathrm{id}(w_{t})]\mathrm{,}$ (1)
where $\bm{W}_{U}\in\mathbb{R}^{|\mathcal{V}|\times d}$ is an unembedding matrix obtained from LM’s output layer, and $|\mathcal{V}|\in\mathbb{R}$ is model’s vocabulary size. Simply put, the internal representaion $\bm{h}_{l,t}$ is mapped into output vocabulary space by applying $\bm{W}_{U}$ (i.e., skipping subsequent layers: $\bm{h}_{l+1,t},\cdots,\bm{h}_{\mathrm{last},t}$), and next-word probability is obtained in that space. $\mathrm{LayerNorm}(\cdot):\mathbb{R}^{d}\to\mathbb{R}^{d}$ in Eq. 1 is the layer normalization at the last layer, and [$\mathrm{id}(w_{t})$] extracts the probability for $w_{t}$333 If a word is split into multiple subwords, accumulated surprisal is used. See Eq.2 in Kuribayashi et al. (2021). from the probability distribution over $\mathcal{V}$, obtained through the $\mathrm{softmax}(\cdot):\mathbb{R}^{|\mathcal{V}|}\to{[0,1]}^{|\mathcal{V}|}$ function. The obtained probability (Eq. 1) is converted to surprisal $-\log p(w_{t}|\bm{w}_{<t};\bm{h}_{l,t})$, and then used in the regression model to predict human reading data (Section˜3.2).
The second method, tuned-lens, extends logit-lens to handle the potential representational drifts through layers. This technique introduces an additional linear transformation for each layer $l$ to mitigate the mismatch between the representation spaces of the $l$-th layer and the last layer:
$p(w|\bm{w}_{<t};\bm{h}_{l,t})$
$=\mathrm{LogitLens}(\bm{W}_{l}\bm{h}_{l,t}+\bm{b}_{l})[\mathrm{id}(w)]\mathrm{,}$ (2)
where $\bm{W}_{l}\in\mathbb{R}^{d\times d}$ and $\bm{b}_{l}\in\mathbb{R}^{d}$ are additionally trained to align the output of logit-lens with the last layer’s next-word probability distribution on additional LM pretraining data. We use publicly available tuned-lens parameters. Notably, we do not fine-tune any part of the LMs for human data; instead, we observe the emerging correlations between next-word probabilities and human reading measures.
3.2 Psychometric predictive power
We evaluate the ability of surprisal values to predict word-by-word human cognitive responses, such as reading times or physiological signals (Section˜4.1). Following prior work (Wilcox et al. 2023b; Pimentel et al. 2023, inter alia), this is done using linear regression models, motivated by surprisal theory Smith and Levy (2013); Shain et al. (2024), which posits a linear relationship between surprisal and processing cost.
Formally, let $\bm{w}=[w_{1},\dots,w_{n}]^{\top}$ denote the sequence of words in a dataset, and let $\bm{y}=[y_{1},\dots,y_{n}]^{\top}\in\mathbb{R}^{n}$ be the corresponding word-by-word human measurements. We assess the predictive contribution of surprisal values $\bm{s}=[s(w_{1}),\dots,s(w_{n})]^{\top}\in\mathbb{R}^{n}_{\geq 0}$, where surprisal is defined as $s(w_{t}):=-\log p(w_{t}|\bm{w}_{<t})$, using model probabilities as described in Section˜3.1.
To evaluate the added benefit of surprisal over more primitive linguistic factors, we include a baseline feature vector $\bm{b}(w_{t})$ for each word $w_{t}$: [^4] Word length is measured in characters; word frequency is estimated using word_freq Speer (2022). We use a consistent set of baseline features across all datasets and models, with a few exceptions. For N400 data, a baseline amplitude term is added; for Michaelov et al. (2024b)’s EEG data, we additionally include electrode-level random effects.
$\bm{b}(w_{t})=[$ $\mathrm{length}(w_{t}),\mathrm{freq}(w_{t}),\mathrm{length}(w_{t-1}),$
$\mathrm{freq}(w_{t-1}),\mathrm{length}(w_{t-2}),\mathrm{freq}(w_{t-2}),$
$s(w_{t-1}),s(w_{t-2})]^{\top}.$ (3)
We include features of the two preceding words to account for spillover effects — i.e., the processing difficulty of $w_{t-1}$ or $w_{t-2}$ can influence the response to $w_{t}$.
We train two nested linear regression models [^5] Implemented using the statsmodels package Seabold and Perktold (2010). Some existing works include preceding words’ surprisal values $s(w_{t-1}),s(w_{t-2})$ only in the full regression model (not in the reduced one; Eq. 3) when computing $\Delta$LL. We confirmed that this setting variation does not alter the conclusion, and at least in this paper, we adopt the setting to include these features in both full and reduced regression models.: (i) a full model including both surprisal and baseline features; and (ii) a reduced model using only the baseline features. The coefficients are estimated via ordinary least squares. Model fit is quantified by the log-likelihood under a Gaussian noise assumption, and the difference in log-likelihoods between the two models — denoted $\Delta\mathrm{LL}$ — reflects the isolated contribution of surprisal.
A higher $\Delta\mathrm{LL}$ indicates greater predictive power, and we refer to this value as the psychometric predictive power (PPP). The central question of our study is: which LM layer yields surprisal values with the strongest PPP?
Figure 2: Relationships between layer depth (x-axis) and $\Delta$LL (y-axis) for each LM in three datasets: FPGD in DC, SPR in NS, and MAZE in NS. The graphs are separated by model families and data. Each line corresponds to a different model, and one with larger markers and more layers corresponds to the bigger model within each model family. For each model, the best layer is indicated with a red edge line and shading patterns, and the last layer is indicated with a black edge. The graph starts at the first layer, not at the embedding layer. The results are based on logit-lens.
PPP_logit-lens PPP_tuned-lens
Stimuli Measure 0-0.2 0.2-0.4 0.4-0.6 0.6-0.8 0.8-1 0-0.2 0.2-0.4 0.4-0.6 0.6-0.8 0.8-1
DC FPGD Kennedy et al. (2003) 14.13 14.78 14.92 13.38 9.84 17.10 16.32 15.39 13.53 10.49
NS SPR Futrell et al. (2021) 9.85 9.75 8.44 5.68 2.67 8.93 7.03 5.11 3.44 2.33
MAZE Boyce and Levy (2023) 1.18 3.00 5.69 12.06 23.77 9.70 17.56 24.15 32.86 39.63
ZuCO FPGD Hollenstein et al. (2018) 38.10 38.13 35.59 29.82 15.94 30.48 27.16 22.56 17.29 8.77
N400 Hollenstein et al. (2018) 0.07 0.12 0.15 0.18 0.16 0.20 0.32 0.34 0.29 0.18
UCL SPR Frank et al. (2013) 22.88 22.21 19.30 11.45 4.77 15.78 8.92 4.87 2.53 1.27
FPGD Frank et al. (2013) 22.11 23.39 22.83 15.77 6.53 16.28 14.48 11.87 9.47 5.57
N400 Frank et al. (2015) 56.86 38.04 22.77 16.07 22.58 11.31 6.12 16.19 29.49 37.11
Fillers SPR Vasishth et al. (2010) 7.83 10.89 14.39 14.18 14.35 8.60 10.47 11.36 11.86 13.33
FPGD Vasishth et al. (2010) 6.66 5.83 6.48 7.31 10.36 8.94 10.91 12.91 13.81 14.00
MAZE Hahn et al. (2022) 5.39 3.01 5.08 21.97 60.89 9.96 28.27 52.00 73.38 88.64
Michaelov+,2024 N400 Michaelov et al. (2024b) 0.88 1.42 1.91 1.68 0.91 0.95 1.51 1.70 1.38 0.99
Federmeier+,2007 N400 Federmeier et al. (2007) 0.77 3.11 8.59 18.05 25.80 1.49 5.22 13.06 24.48 28.71
W&F,2012 N400 Wlotko and Federmeier (2012) 0.35 0.19 0.10 0.09 0.12 0.51 0.27 0.12 0.05 0.11
Hubbard+,2019 N400 Hubbard et al. (2019) 0.18 0.23 0.23 0.25 0.17 0.11 0.12 0.22 0.36 0.33
S&F,2022 N400 Szewczyk and Federmeier (2022) 0.11 0.15 0.38 0.90 1.40 0.16 0.33 0.77 1.29 1.42
Szewczyk+,2022 N400 Szewczyk et al. (2022) 1.21 2.91 4.43 6.40 8.04 2.12 3.58 5.52 8.10 8.93
@@ 4.1 Human data
We use 15 human reading datasets, listed in Table 1, (we additionally use MECO in Section˜6.5), which include human measurements from various methods: self-paced reading time (SPR), first-pass gaze duration (FPGD), Maze task processing time (MAZE), and electroencephalography (EEG; specifically the N400 component). The datasets share a common format: each word $w_{t}$ is annotated with $\mathrm{Cost}(w_{t})\in\mathbb{R}$ representing the human cognitive load associated with it. Our corpus selection aligns with recent studies Kuribayashi et al. (2024); Michaelov et al. (2024a); de Varda et al. (2024); McCurdy and Hahn (2024). [^6] We applied the same preprocessing as Kuribayashi et al. (2024) (DC, NS), de Varda et al. (2024) (UCL), Hahn et al. (2022) (Fillers), and Michaelov et al. (2024a) (N400). For ZuCO, we only used the naturalistic reading part, and for its N400, we averaged the values at the central electrode between 300-500ms during the first pass over a word.